# ZIA: A Theoretical Framework for Zero-Input AI

### v1

@@ 2.2 Brain-Computer Interfaces (BCIs)
BCIs provide a direct precedent for intent prediction from bio-signals, particularly EEG, a core modality in ZIA. Seminal work by [5] established EEG-based control for cursor movement, achieving binary intent classification (e.g., left vs. right) with accuracies of 80-90% in controlled settings. More recent advances, such as DeepConvNet [6], apply convolutional neural networks to raw EEG, improving motor imagery classification to $\sim$85across four classes. These systems rely on explicit user training and predefined task sets, contrasting with ZIA’s goal of unsupervised, free-form intent detection.
Non-invasive BCIs have also explored event-related potentials (ERPs) for intent, with [7] decoding P300 signals for spelling tasks at $\sim$80accuracy. However, ERP-based methods require stimulus-driven paradigms (e.g., visual cues), limiting their applicability to zero-input contexts where no external trigger exists. Invasive approaches, like Neuralink’s primate studies [8], achieve higher precision via implanted electrodes, but their impracticality for consumer use drives ZIA toward scalable, non-invasive solutions. BCI research excels at signal-specific intent but lacks multi-modal integration and edge deployment, key differentiators for ZIA.
@@ 2.4 Edge AI and Real-Time Systems
Edge computing is critical for ZIA’s low-latency, privacy-preserving goals. [13] survey edge AI, reporting inference times of 50–100 ms for vision models on mobile GPUs, leveraging quantization and pruning. MobileNet [14] optimizes convolutional networks for edge devices, achieving $\sim$70accuracy on ImageNet with $\sim$100 ms latency on smartphones. More recently, EfficientNet [15] scales model depth efficiently, but its complexity exceeds ZIA’s real-time needs without further compression.
@@ 3.1.2 Bio-Signal Model
include heart rate ($h_{t}$, sampled at 1 Hz) and EEG ($e_{t}\in\mathbb{R}^{k}$, $k=4-16$ channels, 256 Hz). Heart rate reflects physiological states (e.g., focus vs. stress), modeled as:
@@ 3.3 Edge Constraints
where $C_{\text{cycle}}$ is cycles per operation, Freq is hardware frequency (e.g., 2-4 GHz), and $T_{\text{io}}$ is I/O overhead ($\sim 10$ ms). Standard transformer attention scales as $O(N^{2})$ for sequence length $N$, necessitating optimization (e.g., $O(N)$ linear attention) to meet $T_{\text{inf}}<100$ ms.
@@
where $E_{\text{op}}$ is energy per operation (e.g., 1-2 nJ). ZIA’s edge focus drives trade-offs in model size, precision, and complexity, detailed in Section 5.5.
@@ 4.1.1 Gaze Signal
[…]
@@ 4.1.2 Bio-Signals
include heart rate ($h_{t}$, 1 Hz) and EEG ($e_{t}\in\mathbb{R}^{k}$, $k=4-16$, 256 Hz). Heart rate, indicative of physiological states like focus or stress, is smoothed:
@@ 5.1.2 Multi-Modal Contributions
Adding heart rate ($s_{b}(t)_{h}$) increases $I(I_{t};s_{g},s_{b})$ by $0.3-0.5$ bits, reflecting physiological cues (e.g., focus elevates HR). Contextual data ($s_{c}(t)$)—time, location, and usage history—adds $0.5-0.7$ bits via situational priors. EEG ($s_{b}(t)_{e}$), with $k=16$ channels and SNR $\approx 10$ dB, contributes $0.8-1.0$ bits due to neural intent correlates (e.g., beta band activity). Cumulatively:
@@ 5.2.1 Standard Transformer
For an Edge TPU (4 GHz, $C_{\text{cycle}}\approx 1$), latency reduces to $60-70$ ms, and on an NPU (3 GHz), it is $70-80$ ms.
@@ 5.2.2 Linear Attention Variant
[…]
TPU achieves $35-45$ ms, and NPU achieves $40-50$ ms—both well under 100 ms.

### v3

@@ 2.2 Brain-Computer Interfaces (BCIs)
BCIs provide a direct precedent for intent prediction from bio-signals, particularly EEG, a core modality in ZIA. Seminal work by [5] established EEG-based control for cursor movement, achieving binary intent classification (e.g., left vs. right) with accuracies of 80-90% in controlled settings. More recent advances, such as DeepConvNet [6], apply convolutional neural networks to raw EEG, improving motor imagery classification to $\sim$85% across four classes. These systems rely on explicit user training and predefined task sets, contrasting with ZIA’s goal of unsupervised, free-form intent detection.
Non-invasive BCIs have also explored event-related potentials (ERPs) for intent, with [7] decoding P300 signals for spelling tasks at $\sim$80% accuracy. However, ERP-based methods require stimulus-driven paradigms (e.g., visual cues), limiting their applicability to zero-input contexts where no external trigger exists. Invasive approaches, like Neuralink’s primate studies [8], achieve higher precision via implanted electrodes, but their impracticality for consumer use drives ZIA toward scalable, non-invasive solutions. BCI research excels at signal-specific intent but lacks multi-modal integration and edge deployment, key differentiators for ZIA.
@@ 2.4 Edge AI and Real-Time Systems
Edge computing is critical for ZIA’s low-latency, privacy-preserving goals. [13] survey edge AI, reporting inference times of 50–100 ms for vision models on mobile GPUs, leveraging quantization and pruning. MobileNet [14] optimizes convolutional networks for edge devices, achieving $\sim$70% accuracy on ImageNet with $\sim$100 ms latency on smartphones. More recently, EfficientNet [15] scales model depth efficiently, but its complexity exceeds ZIA’s real-time needs without further compression.
@@ 3.1.2 Bio-Signal Model
include heart rate ($h_{t}$, sampled at 1 Hz) and EEG ($e_{t}\in\mathbb{R}^{k}$, k=416k=4-16 channels, 256 Hz). Heart rate reflects physiological states (e.g., focus vs. stress), modeled as:
@@ 3.3 Edge Constraints
where $C_{\text{cycle}}$ is cycles per operation, Freq is hardware frequency (e.g., 2-4 GHz), and $T_{\text{io}}$ is I/O overhead ($\sim 10$ ms). Standard transformer attention scales as $O(N^{2})$ for sequence length NN, necessitating optimization (e.g., $O(N)$ linear attention) to meet $T_{\text{inf}}<100$ ms.
@@
where $E_{\text{op}}$ is energy per operation (e.g., 1-2 nJ). ZIA’s edge focus drives trade-offs in model size, precision, and complexity, detailed in Section $5.$
@@ 4.1.1 Gaze Signal
$S_{g}(f)=\frac{2b_{g}^{2}}{b_{g}^{2}+(2\pi f)^{2}}\quad\text{(Laplacian)}\quad\text{vs.}\quad S_{g}(f)=\frac{\sigma_{g}^{2}}{1+(2\pi f\tau)^{2}}\quad\text{(Gaussian)}$
@@ 4.1.2 Bio-Signals
include heart rate ($h_{t}$, 1 Hz) and EEG ($e_{t}\in\mathbb{R}^{k}$, k=416k=4-16, 256 Hz). Heart rate, indicative of physiological states like focus or stress, is smoothed:
@@ 5.1.2 Multi-Modal Contributions
Adding heart rate ($s_{b}(t)_{h}$) increases $I(I_{t};s_{g},s_{b})$ by 0.30.50.3-0.5 bits, reflecting physiological cues (e.g., focus elevates HR). Contextual data ($s_{c}(t)$)—time, location, and usage history—adds 0.50.70.5-0.7 bits via situational priors. EEG ($s_{b}(t)_{e}$), with $k=16$ channels and SNR $\approx 10$ dB, contributes 0.81.00.8-1.0 bits due to neural intent correlates (e.g., beta band activity). Cumulatively:
@@ 5.2.1 Standard Transformer
For an Edge TPU (4 GHz, $C_{\text{cycle}}\approx 1$), latency reduces to 607060-70 ms, and on an NPU (3 GHz), it is 708070-80 ms.
@@ 5.2.2 Linear Attention Variant
$T_{\text{inf}}=\frac{5\times 10^{7}\cdot 2}{2.84\times 10^{9}}+10\approx 45-55\,\text{ms}\quad\text{(CPU)}.$
TPU achieves 354535-45 ms, and NPU achieves 405040-50 ms—both well under 100 ms.